OpenStudy (anonymous):

\(\sf \Large \hspace{50 pt} \text{Discriminant of Quadratic Function}\\ \sf \hspace{100 pt} \text{by Data_LG2}\) In a quadratic function (\(\sf ax^2+bx+c=0\)), the number or types of roots it will have can be determined by its \(\sf discriminant\). This topic will be discussed below. Source: My Peabrain.

OpenStudy (anonymous):

\(\sf \Large \color{red}{Discriminant}\) came from the famous \(\sf \text{quadratic formula}: \\ \hspace{-75 pt} \Large x=\frac{{-b \pm \sqrt{\color{blue}{b^2-4ac}}}}{2a}\) The discriminant is given as: \(\sf \large \color{red}{D}= \color{blue}{b^2-4ac}\) **Notice that this formula can be seen in the radical from the quadratic formula. This determines the types of roots that the function will have depending on the following conditions: \(\hspace{20 pt} \begin{array}{|l|c|} \hline \sf \large \text{If D is...}&\sf \large \text{then quadratic function will have...}\\ \hline \hspace{5 pt} \sf < 0 &\sf \text{no real roots}\\ \hline \hspace{5 pt} \sf = 0 &\sf \text{one real root}\\ \hline \hspace{5 pt} \sf > 0 &\sf \text{two real roots}\\ \hline \end{array}\)

OpenStudy (anonymous):

\(\sf \Large \text{Equation is Given}\) 1. Make sure that the equation is in the standard form \(\sf ax^2 + bx +c =0\), otherwise rearrange the equation. 2. Identify the values of a, b, and c from the equation. 3. Plug-in the coefficient values into the discriminant formula to solve for the discriminant. From here, you can determine the type of roots the quadratic function will have.

OpenStudy (anonymous):

\(\sf \huge \color{blue}{\text{Sample Questions:}}\\ \sf \text{How many solutions will the following quadratic functions will have?} \) A. \(\sf \color{#999900}{-3x^2 + 2x + 8 = 0}\) 1. The equation is in std form. 2. Values for a, b, and c are -3, 2, and 8 respectively. 3. Plug it into the formula: \(\sf b^2-4ac\\ \sf (2)^2-4(-3)(8) \\ \sf = 100\) Therefore, since the value of the discriminant is 100, which is greater than zero and it is a perfect square of 10, the quadratic function will have \(\sf \text{Two Rational Solutions}\) B. \(\sf \color{#999900}{2x^2 - 5x = -20} \) 1. The equation is not in std form, so let's rearrange it a little bit. \(\hspace{15 pt} \sf 2x^2 - 5x \color{red}{+ 20} = -20 \color{red}{+ 20} \\ \hspace{15 pt} \sf 2x^2 -5x + 20 = 0\) 2. Values for a, b, and c are 2, -5, and 20 respectively. 3. Plug it into the formula: \(\sf b^2-4ac\\ \sf (-5)^2-4(2)(20) \\ \sf = -135\) Since the discrimant is -135, which is less than zero, this quadratic function will have \(\sf \text{No Real Solutions}\) (imaginary). C. \(\sf \color{#999900}{2x^2 - 4x + 2 = 0}\) 1. 1. The equation is in std form. 2. Values for a, b, and c are 1, -6, and 13 respectively. 3. Plug it into the formula: \(\sf b^2-4ac\\ \sf (4)^2-4(2)(2) \\ \sf = 0\) The discriminant is equal to zero, this means that this quadratic function has \(\sf \text{One Real Solution}\).

OpenStudy (anonymous):

[End]

OpenStudy (abhisar):

That's nice and clean, thank you for sharing with us (-:

OpenStudy (anonymous):

Oh wow. Very nice explanation

Nnesha (nnesha):

veryyyyy nice!!!